18 cm Session

Session: Feb 2002 (session 1)
Wavelength: 18 cm


To give feedback on this session, click here

18 cm Session Feedback

Torun (Kaz Borkowski at Mon Feb 11 10:02:35 MET 2002) Minor failures
1) One of our 8 baseband converters (BBC 7) is out of order. This will affect all experiments using 8 BBCs. 2) Due to a problem with the I.F. Distributor, Field System cannot carry out the Tsys measurements. To amend this loss the operators were asked to make regular Tsys measurements manually. These manual measurements however are done in wide bandwitdth.

Jodrell_Bank (Alastair Gunn at Tue Feb 12 11:24:20 MET 2002) Minor failures
It appears that tpi (and therefore Tsys measurements) in our VC12 device are corrupt. Will investigate problem after session. Please note that in experiment EC018 we recorded the Mk2 telescope for the first three tapes, then switched to Lovell telescope for final three.

Cambridge (Alastair Gunn at Tue Feb 12 11:26:38 MET 2002) Failure
At the present time the Cambridge telescope is unable to observe. Orange, the mobile phone company, have damaged one of the L-band microwave repeaters between Cambridge and Jodrell Bank. Time to fix is currently unknown.

Westerbork (Tony Foley at Mon Feb 25 14:10:08 MET 2002) Minor failures
Most of the session I used regular TPI measurements on BBCs 1-8 (irrespective of which were used) as I could then automate plotting with LOGPL. The measurements were done every 13 seconds so that they would rarely beat with the 10second noise switching period. The phasing of X and Y linear to RCP and LCP was changed on day 54 after feedback from the NME.